Use the given conditions to write an equation for the line in​ point-slope form and in​ slope-intercept form.
Passing through ​(3​,-2​) and perpendicular to the line whose equation is y=1/2x+4. What's the point-slope form
1
Expert's answer
2020-09-14T18:48:45-0400

The products of the gradients of perpendicular lines is "-1." That is to say, "mm' = -1"


Let's assume that the gradient of the line perpendicular to the line "y = \dfrac {1} {2} x + 4" is "m."

"=> m(\dfrac {1}{2}) = -1"

"=> \dfrac {1}{2}m = -1"

"=> m = -2"


Point-slope form

"ANSWER: \space y + 2 = -2(x - 3)"


"SOLUTION"

The general point-slope form of a linear equation is

"y - y_{1} = m(x - x_{1})" where "m" is the slope/gradient and "(x_{1}; y_{1})" is the point.

In this case, "m = -2 \space and \space points \space x_{1} = 3 \space \& \space y_{1} = -2 \space from \space the \space point \space (3; -2)"


"\therefore \space y - (-2) = -2(x-3)"

"=> y + 2 = -2 (x - 3)"



Slope-intercept form

"ANSWER: y = -2x + 4"


"SOLUTION"

The general slope-intercept form of a linear equation is given by "y = mx + c" where "m" is the slope/gradient and "c" is the "y-intercept."

Now, "m = -2" , therefore "y = -2x + c"

Substituting the point "(3; -2)" to find "c:" "-2 = -2(3) + c"

"=> -2 = -6 + c"

"=> c = -2 + 6"

"=> c = 4"

"\therefore \space y = -2x + 4"